ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

6. Topology

1991 MS Classification: 57 Morton, Hugh R. University of Liverpool, England The Burau matrix and Fiedler’s invariant for closed braids I will show how Fiedler’s ‘small state-sum’ invariant for a closed braid β̂ can be found in terms of the Burau representation of β, and hence from the 2-variable Alexander polynomial of the link β̂ ∪ A consisting of the closed braid β̂ and its axis A.